### v1

# Universality of Dissipation across Holographic Interfaces

Abstract
Motivated by recent results in spin chains we study dissipation and relaxation in a two-dimensional holographic interface conformal field theory (ICFT) in which degrees of freedom on one side of the interface are coupled to an external bath, while the other side remains isolated. In the bulk description this setup is realized by gluing a supersymmetric Janus geometry to a BTZ black hole region, with the coupling implemented through a double-trace deformation. We determine the quasinormal modes in the bulk by solving the double-trace matching conditions of the system and bath. The lowest imaginary part of the modes defines a Liouvillian gap, and following earlier work in spin chains we introduce the dimensionless ratio $c_{\rm relax}$ as a measure of interface-induced suppression of relaxation. Numerically we find that, $c_{\rm relax}$ is independent of coupling details to the bath. It is a strong candidate for a universal interface observable characterizing dissipation and relaxation across the interface.
1 Introduction
Conformal interfaces in two-dimensional quantum field theories provide a rich arena for probing universal properties of strongly coupled systems. In two-dimensional interface conformal field theories (ICFTs), an interface separates two conformal fixed points while preserving a diagonal conformal symmetry. Despite their local nature, such interfaces are characterized by a small set of universal quantities that capture intrinsic properties independent of microscopic details. Well-known examples include the effective central charge $c_{\text{eff}}$, the energy transmission coefficient $c_{LR}$, and the boundary entropy or g-factor, all of which admit precise definitions within field theory and have elegant realizations in holographic duals Affleck and Ludwig (1991); Quella et al. (2007); Meineri et al. (2020); Karch and Wang (2023); Karch et al. (2023); Chiodaroli et al. (2010a, 2017).
From the bulk perspective, these universal ICFT data are naturally realized in Janus geometries, where the bulk scalar profile interpolates between distinct asymptotic AdS regions. In this framework, the interface observables can be extracted geometrically from minimal surfaces calculating entanglement entropies, or from a bulk mode analysis. They are insensitive to continuous deformations that preserve the interface conformal structure. Supersymmetric extensions, including super-Janus solutions, further enrich this picture by allowing controlled embeddings into string theory and enabling the study of supersymmetric ICFTs with exact control over protected quantities.
While much of the existing literature has focused on static or equilibrium characteristics of conformal interfaces, considerably less is known about their dynamical and dissipative properties. In particular, when an interface system is coupled to an external environment, new universal structures may emerge that are not visible in equilibrium observables. Understanding such dissipative dynamics is of both conceptual and practical interest, as it connects interface CFTs to broader questions of open quantum systems, relaxation, and information loss at strong coupling.
Recently, progress in this direction was made in Barad et al. (2025), where dissipation at conformal interfaces was studied using a combination of field-theoretic and holographic techniques. A central insight of that work is that the late-time relaxation of interface degrees of freedom can be characterized by a universal dissipation rate governed by a Liouvillian gap Baumgartner and Narnhofer (2008). This suggests the existence of new interface data, $c_{\text{relax}}$, controlling nonequilibrium dynamics, analogous in spirit to $c_{\text{eff}}$, $c_{LR}$, and $g$, but intrinsically dynamical in nature.
Motivated by these developments, in this paper we approach the problem holographically and propose a universal quantity characterizing dissipation at supersymmetric Janus interfaces, which we also denote by $c_{\text{relax}}$. Like in the spin-chain example of Barad et al. (2025) this quantity is defined through the ratio of lowest relaxation rates of scalar modes in the bulk, where a BPS-Janus supergravity Chiodaroli et al. (2010b, 2017) is coupled to a thermal BTZ bath via double-trace deformation on one side of the interface. It is dual to a dissipative observable in a two-dimensional supersymmetric ICFT.
Our main result is that the resulting relaxation ratio $c_{\text{relax}}$ exhibits a striking universality similar to the one observed in spin chains. Namely, the ratio of the interface relaxation rate to that of a planar AdS3 system is independent of the coupling strength, the microscopic details of the coupling characterized by a dimension $\Delta$ and, in setups with localized coupling, independent of the position along the interface. This universality mirrors the behavior of equilibrium ICFT data such as $c_{\text{eff}}$ and $c_{LR}$, and strengthens the case for $c_{\text{relax}}$ as a new universal characteristic of dissipative interface dynamics in two-dimensional supersymmetric ICFTs.
This paper is organized as follows. In Section 2, 3 and 4 we review the normal and quasinormal modes in pure AdS3, BPS-Janus and BTZ black hole, respectively. In Section 5 we formulate the double-trace deformation equations between the coupling of BPS-Janus and BTZ in two ways: half-space coupling and a localized single-point coupling. In Section 6 we present the numerical results and demonstrate the universality of $c_{\text{relax}}$. We also compare it with known equilibrium interface observables like $c_{\text{eff}}$ and $c_{LR}$ in BPS-Janus. Finally in Section 7 we discuss future directions.
2 Normal modes of empty AdS3
In this section, we review the normal modes of empty AdS3 without an interface. Although the results are standard, in order to map the results to Janus geometry, we approach it in AdS2 slicing coordinates rather than the usual cylinder picture.
Consider the foliation of asymptotic AdS3 by AdS2 slices
$ds^{2}=dx^{2}+e^{2A(x)}\,\sec^{2}\sigma\,(-dt^{2}+d\sigma^{2})$ (1)
$x\to\pm\infty$ are boundary pieces from left and right, and $\sigma\in[-\tfrac{\pi}{2},\tfrac{\pi}{2}]$. For empty AdS3 we have $e^{A}=\cosh x$. Throughout the paper the AdS3 radius is set to $\ell=1$. A probe scalar with mass $M$ in AdS3 solved by the Klein–Gordon (KG) equation can be decomposed as
$\Psi_{\omega}=\sum_{n}f_{n}(x)g_{\omega,n}(\sigma)e^{-i\omega t}$ (2)
where $g_{\omega,n}$ are the eigenfunctions of the AdS2 radial KG equation with mass $m_{n}$, and $f_{n}(x)$ solves the slicing modes equation. Following results of Aharony et al. (2003); Karch and Sato (2017), for general warpfactor $e^{A(x)}$ the equation for the slicing direction is
$f^{\prime\prime}(x)+2A^{\prime}(x)f^{\prime}(x)=(M(x)^{2}-e^{-2A(x)}m^{2})f(x)$ (3)
For empty AdS3, the mass tower is
$\Delta_{n}=\Delta+n,\quad n=0,1,\dots$ (4)
where $m_{n}^{2}=\Delta_{n}(\Delta_{n}-1)\equiv\nu_{n}(\nu_{n}+1)$ and $\Delta\equiv\Delta_{+}=1+\sqrt{M^{2}+1}$. The functions $f$ and $g$ can be written out explicitly Karch and Sato (2017); Higuchi et al. (2022); Spradlin and Strominger (1999)
[…]
Note that the two linearly-independent solutions for $g$ have different parities across the spatial infinity at two boundary points $\sigma=\pm\tfrac{\pi}{2}$.
Near the boundary of AdS3, the normalizable vev $\beta$ and non-normalizable source $\alpha$ defined by
$\Psi_{\omega}\sim\alpha_{\omega}(\sigma)\,e^{-(d-\Delta)|x|}+\beta_{\omega}(\sigma)\,e^{-\Delta|x|}$ (6)
are read off by the AdS2 modes near the expansion, via the BOPE formula Aharony et al. (2003)
$\mathcal{O}_{d=2}(\sigma)=\sum_{n}\frac{b_{\Delta_{\mathcal{O}}\Delta_{n}}}{\tan(\tfrac{\sigma}{2}+\tfrac{\pi}{4})^{\Delta_{\mathcal{O}}-\Delta_{n}}}\mathcal{O}_{d-1,n}$ (7)
We can derive the AdS3 normal modes by picking out the normalizable solution near the boundary, while demanding it to vanish near infinity Birmingham (2001). Transforming the hypergeometric function gives Higuchi et al. (2022)
$\begin{split}g_{\omega,n}^{(1)}(\sigma)=(\cos\sigma)^{\Delta_{n}}\,A(\omega)\,F\!\left(\frac{\Delta_{n}+\omega}{2},\,\frac{\Delta_{n}-\omega}{2};\,\frac{1}{2}+\Delta_{n};\,\cos^{2}\sigma\right)\\ +(\cos\sigma)^{1-\Delta_{n}}\,B(\omega)\,F\!\left(\frac{1-\Delta_{n}+\omega}{2},\,\frac{1-\Delta_{n}-\omega}{2};\,\frac{3}{2}-\Delta_{n};\,\cos^{2}\sigma\right).\end{split}$ (8)
where
$A_{1}(\omega):=\frac{\Gamma\!\left(\tfrac{1}{2}\right)\,\Gamma\!\left(\tfrac{1}{2}-\Delta_{n}\right)}{\Gamma\!\left(\tfrac{1-\Delta_{n}+\omega}{2}\right)\,\Gamma\!\left(\tfrac{1-\Delta_{n}-\omega}{2}\right)},\qquad B_{1}(\omega):=\frac{\Gamma\!\left(\tfrac{1}{2}\right)\,\Gamma\!\left(\Delta_{n}-\tfrac{1}{2}\right)}{\Gamma\!\left(\tfrac{\Delta_{n}+\omega}{2}\right)\,\Gamma\!\left(\tfrac{\Delta_{n}-\omega}{2}\right)}.$ (9)
We perform the same transformation for $g^{(2)}$ giving $A_{2},B_{2}$. The vanishing condition demands $B_{1,2}(\omega)=0$ which combining the two even/odd parities gives the discrete empty AdS3 normal modes
$\omega_{k,n}=\Delta_{n}+k=\Delta+n+k,\quad k,n\in\mathbb{N}_{\geq 0}$ (10)
This matches the global AdS3 results previously obtained using cylinder coordinates Spradlin and Strominger (1999). At these discrete values $\omega=\omega_{k,n}$, the Legendre functions in $g$ become Jacobi polynomials that form an orthonormal basis (here with $\nu=\nu_{n}=\Delta_{n}-1$)
$g_{\nu,k}(\sigma)=N_{\nu,k}\,(\cos\sigma)^{\nu+1}\,P^{(\nu+\tfrac{1}{2},\,\nu+\tfrac{1}{2})}_{k}(\sin\sigma),$ (11)
where
$P^{(a,b)}_{m}(x):=\frac{\Gamma(m+a+1)}{m!\,\Gamma(a+1)}\,{}_{2}F_{1}\!\left(-m,\,m+a+b+1;\,a+1;\,\tfrac{1-x}{2}\right).$ (12)
and
$N_{\nu,k}=\frac{\sqrt{k!\,\Gamma(2\nu+2+k)}}{2^{\nu+1}\Gamma(\nu+k+\tfrac{3}{2})}$ (13)
3 Normal modes of BPS-Janus
[…]
The ten-dimensional metric can be written as a fibration over a Riemann surface $\Sigma$ with boundary,
$ds^{2}_{10}=f_{1,10}^{2}\,ds^{2}_{AdS_{2}}+f_{2,10}^{2}\,ds^{2}_{S^{2}}+\rho_{10}^{2}\,dw\,d\bar{w}+f_{3,10}^{2}\,ds^{2}_{T^{4}},\qquad w\in\Sigma,$ (14)
with all warp factors depending only on $w,\bar{w}$. It is convenient to reduce to six dimensions (Einstein frame) and write
$ds_{6}^{2}=f_{1}^{2}\,ds^{2}_{AdS_{2}}+f_{2}^{2}\,ds^{2}_{S^{2}}+\rho^{2}\,dw\,d\bar{w},$ (15)
together with identities
$f_{1}^{2}f_{2}^{2}=H^{2},\qquad\sqrt{-g}=\rho^{2}H^{2},$ (16)
where $H$ is one of the harmonic functions specifying the solution.
We focus on the simplest Janus solutions with two asymptotic AdS3 regions and $\Sigma$ an infinite strip. Writing
$w=x+iy,\qquad x\in\mathbb{R},\quad y\in(0,\pi),$ (17)
the two asymptotic AdS3 regions are approached as $x\to\pm\infty$. Nontrivial Janus deformations are parameterized by two real parameters $(\psi,\theta)$ controlling, respectively, the dilaton and axion jumps; the undeformed limit is $\psi=\theta=0$. A convenient combination is the interface parameter
$\kappa\equiv\cosh\psi\,\cosh\theta.$ (18)
For the RR-supported Janus family, the harmonic function and the ratios of metric factors take the simple form
$H=2\,\cosh x\,\sin y,\qquad\frac{\rho^{2}}{f_{1}^{2}}=\frac{\kappa^{2}}{\cosh^{2}x},\qquad\frac{\rho^{2}}{f_{2}^{2}}=\frac{1}{\sin^{2}y}+\frac{\kappa^{2}-1}{\cosh^{2}x}.$ (19)
Let us consider a scalar field $\Phi$ dual to a CFT operator of scaling dimension $\Delta$, propagating in the six-dimensional BPS Janus background. In the asymptotic AdS3 regions, the bulk mass is related to $\Delta$ by $m^{2}=\Delta(\Delta-2)$.
The Klein–Gordon (KG) equation may be written as
$\left(\Box_{6}-m^{2}\right)\Phi=0,$ (20)
or, using the metric (15) and the identity (16),
$\frac{1}{\rho^{2}}\left[\frac{\partial_{a}\!\left(H^{2}\partial_{a}\right)}{H^{2}}+\frac{\rho^{2}}{f_{1}^{2}}\nabla^{2}_{AdS_{2}}+\frac{\rho^{2}}{f_{2}^{2}}\nabla^{2}_{S^{2}}\right]\Phi-\Delta(\Delta-2)\,\Phi=0,\qquad a=x,y.$ (21)
Despite the presence of the mass term, the equation remains fully separable
$\Psi_{\omega}=f_{Ln}(x)g_{knLl}(\sigma)e^{-i\omega t}Y_{Ll}(y,\theta,\phi)$ (22)
with
$\nabla^{2}_{AdS_{2}}f=\hat{\nu}(\hat{\nu}+1)f,\qquad\nabla^{2}_{S^{2}}Y=-l(l+1)Y.$ (23)
Using the explicit Janus metric ratios (19), the remaining $(x,y)$ equation becomes
$\left[\partial_{x}\!\left(\cosh^{2}x\,\partial_{x}\right)+\partial_{y}\!\left(\sin^{2}y\,\partial_{y}\right)-\frac{l(l+1)}{\sin^{2}y}+\frac{\nu(\nu+1)}{\cosh^{2}x}-\Delta(\Delta-2)\,H^{2}\right]\Psi_{\hat{\nu}l}(x,y)=0,$ (24)
and regularity demands the quantization of parameters
$\nu_{n}(\nu_{n}+1)\equiv(\Delta+n-1)(\Delta+n)=l(l+1)+\kappa^{2}\!\left[\hat{\nu}_{n}(\hat{\nu}_{n}+1)-l(l+1)\right].$ (25)
[…]
$\omega_{m,n,L,l}=\hat{\Delta}_{n}+k$ (26)
where
$\hat{\Delta}_{n}=\left\{\frac{1}{4}+\frac{(L+n+\Delta-1)(L+n+\Delta)}{\kappa^{2}}+l(l+1)\left(1-\frac{1}{\kappa^{2}}\right)\right\}^{1/2}+\frac{1}{2}$ (27)
If we demand $L=l=0$ and only consider the quantum numbers in the asymptotic AdS3 then the tower of normal modes is
$\omega_{n,k}=k+\frac{1}{2}+\left(\frac{1}{4}+\frac{(n+\Delta-1)(n+\Delta)}{\kappa^{2}}\right)^{1/2},\quad n,k\in\mathbb{N}_{\geq 0}$ (28)
4 Quasinormal modes in Schwarzschild BTZ black hole
The thermal bath that we are coupling our BPS-Janus system to is the BTZ black hole on the same geometry, on both sides of the interface. In this section we review the quasinormal modes of the BTZ black hole Birmingham (2001); Cardoso and Lemos (2001).
If we use the conventional cylinder coordinates for BTZ with $\ell=1$, $J=0$,
$ds^{2}=-(r^{2}-r_{+}^{2})\,dt^{2}+\frac{1}{r^{2}-r_{+}^{2}}\,dr^{2}+r^{2}\,d\phi^{2},\qquad\phi\sim\phi+2\pi.$ (29)
The KG equation in BTZ for a scalar with dimension $\Delta_{BTZ}$ has solution $\Psi^{BTZ}=R(r)e^{-i\omega t}e^{im\phi}$. Solving the radial equation gives
$R(r)=z^{-i\tfrac{\omega}{2r_{+}}}(1-z)^{\tfrac{2-\Delta_{BTZ}}{2}}\,_{2}F_{1}(a,b;c;z)$ (30)
with $z=1-\tfrac{r_{+}^{2}}{r^{2}}$, and $a,b=\tfrac{2-\Delta_{BTZ}}{2}+i\tfrac{1}{2r_{+}}(-\omega\pm m)$, $c=1-i\tfrac{\omega}{r_{+}}$. Near the boundary $z\to 1$ or $r\to\infty$, we have expansion
$\Psi_{\omega}^{BTZ}\sim\sum_{m}(\alpha_{m}^{BTZ}\,r^{-(2-\Delta_{BTZ})}+\beta_{m}^{BTZ}\,r^{-\Delta_{BTZ}})e^{im\phi}$ (31)
The normalizable and non-normalizable modes are
$\begin{split}\alpha_{m}^{BTZ}(\omega,\Delta_{BTZ})\;=\;\frac{\Gamma(\Delta_{BTZ})\,\Gamma\!\left(1-\tfrac{i\omega}{r_{+}}\right)}{(\Delta_{BTZ}-1)\,\Gamma\!\left(\tfrac{\Delta_{BTZ}-i(\omega-m)/r_{+}}{2}\right)\,\Gamma\!\left(\tfrac{\Delta_{BTZ}-i(m+\omega)/r_{+}}{2}\right)}\\ \beta_{m}^{BTZ}(\omega,\Delta_{BTZ})\;=\;\frac{\Gamma(1-\Delta_{BTZ})\Gamma\!\left(1-\tfrac{i\omega}{r_{+}}\right)}{\Gamma\!\left(\tfrac{-\Delta_{BTZ}-i(\omega-m)/r_{+}+2}{2}\right)\,\Gamma\!\left(\tfrac{-\Delta_{BTZ}-i(m+\omega)/r_{+}+2}{2}\right)}\end{split}$ (32)
The quasinormal modes of the BTZ black hole without rotation is then (after demanding that $\alpha$ vanishes) Cardoso and Lemos (2001)
$\omega=\pm m-ir_{+}(\Delta_{BTZ}+2n)$ (33)
5 Coupling BPS-Janus to BTZ bath
In this section we introduce dissipation to the interface CFT by coupling it to an external thermal bath CFT via a double-trace deformation. From the bulk point of view, we consider two separate spacetimes, one being the Janus geometry and the other a standard BTZ black hole, and let them communicate via boundary conditions along parts of their boundary. From the CFT perspective this corresponds to coupling part of the ICFT to a thermal bath. Both the zero temperature ICFT and finite temperature bath live on the same line, but they are completely decoupled on one side of the interface and only exchange energy on the other side.
The coupling of the interface CFT and the bath is through a double-trace (DT) deformation where one considers two initially decoupled CFTs, a “system” and a “bath”, interacting through a marginal deformation of the form Aharony et al. (2001); Witten (2002); Aharony et al. (2006); Kiritsis (2006); Giombi et al. (2025); Karch et al. (2024b); Hartman and Rastelli (2008)
$S_{\text{int}}=h\int d^{d}x\,\mathcal{O}_{\text{sys}}(x)\,\mathcal{O}_{\text{bath}}(x),$ (34)
with $\Delta_{\text{sys}}+\Delta_{\text{bath}}=d$. Below our system will be the Janus ICFT, so $\Delta_{\text{sys}}=\Delta_{J}$, and the bath the BTZ black hole, so $\Delta_{\text{bath}}=\Delta_{BTZ}$. In the bulk, this interaction is implemented by transparent boundary conditions relating the normalizable and non-normalizable modes of dual bulk fields across two asymptotically AdS spacetimes. Although the combined system evolves unitarily, the subsystem corresponding to the original CFT exhibits dissipative behavior, manifested through energy fluxes and the appearance of quasinormal modes with nonzero imaginary frequencies.
There are two natural ways to implement the DT deformation on one side of the interface. One way is to turn on the DT deformation on the entire half space of the boundary. Another way is to only turn on the deformation at one point, corresponding to a delta-function source.
In the following we solve for the quasinormal modes of the BPS-Janus system coupled to BTZ from the DT equations in both scenarios. The recipe is as follows: We first relax the normalizable boundary condition and write out the modes $\alpha_{\omega}$ and $\beta_{\omega}$ to solve the DT deformation equations between BPS-Janus and temperature-$T$ BTZ modes. To do so, we use the orthogonal basis on the AdS2 slice to decompose the modes. For BPS-Janus, demanding the momenta $k,l$ on $S^{3}$ being zero (integrating out the internal KK modes), the radial solution $f(x)$ is the same with empty AdS3. From the BOPE equation (7), we can decompose the boundary modes on the CFT2 into boundary modes of the AdS2 slice. That means for this case, we can solve the DT deformation equations regarding $\alpha(\sigma),\beta(\sigma)$ on AdS2 slice level, near the AdS2 boundary.
It is worth mentioning that for bottom-up Janus without supersymmetry, we will have to go the other way: First derive $\alpha,\beta$ from the radial fluctuation solution relaxing the infinity regularization condition. Then decompose $\alpha_{\omega}(\sigma)=\sum_{m,n}\alpha_{\omega}g_{m,n}(\sigma)$ and transform between Janus basis and BTZ/empty AdS3 basis to solve the DT mode equations. Following the Heun solutions and the BOPE tower in Gutperle and Yeo (2025) we can solve for the quasinormal modes order by order in the Janus parameter. We will leave this for future work.
5.1 Scalar fluctuations in BPS-Janus
Let us first write down the solution explicitly for the KG equations in BPS-Janus with $L=l=0$. From Chiodaroli et al. (2017), the radial function $f(x)$ in (22) is the same as the undeformed case in (5). For the KG equation of a scalar of dimension $\Delta_{J}$, The general solution with AdS2 BOPE dimension $\nu+1$ that is regular in the bulk is
$f_{\nu}(x)=\frac{C_{\nu}\Gamma(\Delta_{J})}{2\cosh x}P_{-\nu-1}^{1-\Delta_{J}}(\tanh x),\qquad C_{\nu}^{2}=\frac{2\,\Gamma(\nu+\Delta_{J})(2\nu+1)}{\Gamma(\nu-\Delta_{J}+2)\,\Gamma(\Delta_{J})^{2}}$ (35)
In writing this solution we assumed the normalizability of the right side boundary, while engaging the double-trace deformation coupling between BTZ and Janus on the left side.
On the other hand, we do not wish to engage the DT deformation on the interface, which means that on each AdS2 slice, we still impose the normalizability condition, which forces the solution $g_{\omega,\hat{\nu}}(\sigma)$ in the form of (2), with eigenvalues $\omega=\hat{\nu}+1+k$, for $k=0,1,\dots$. Here $\nu(\nu+1)=\kappa^{2}\hat{\nu}(\hat{\nu}+1)$. Together, the eigenfunction under fixed $\omega$ is
$\Psi^{Janus}_{\omega}=\sum_{k\geq 0}f_{\nu(k)}(x)g_{\omega,k}(\sigma)$ (36)
with $\nu(k)\left(\nu(k)+1\right)=\kappa^{2}(\omega-k-1)(\omega-k)$, and
$g_{\omega,k}(\sigma)=\frac{\sqrt{k!\,\Gamma(2\omega-k)}}{2^{\omega-k}\Gamma(\omega+\tfrac{1}{2})}\,(\cos\sigma)^{\omega-k}\,P^{(\omega-k-\tfrac{1}{2},\,\omega-k-\tfrac{1}{2})}_{k}(\sin\sigma)$ (37)
Note that the integer $k$ is bounded $0\leq k\leq\mathrm{Re}(\omega)$ for the solution to be regular near the boundaries $\sigma=\pm\tfrac{\pi}{2}$. Near one side of the boundary hemisphere $x\to$-$\infty$ where we introduce the DT coupling, we have the normalizable and non-normalizable modes expanded in the basis of $g_{\omega,k}$:
$\Psi_{\omega}^{Janus}\sim\sum_{k}\left(\alpha_{k}\,g_{\omega,k}(\sigma)e^{(2-\Delta_{J})x}+\beta_{k}\,g_{\omega,k}(\sigma)e^{\Delta_{J}\,x}\right)$ (38)
where
$\alpha_{k}^{J}(\omega,\Delta_{J})=\frac{C_{\nu(k)}\Gamma(\Delta_{J})^{2}}{(\Delta_{J}-1)\Gamma(\Delta_{J}-\nu(k)-1)\Gamma(\Delta_{J}+\nu(k))},\qquad\beta_{k}^{J}(\omega,\Delta_{J})=-\frac{C_{\nu(k)}\Gamma(\Delta_{J})\Gamma(1-\Delta_{J})}{\Gamma(\nu(k))\Gamma(1-\nu(k))}$ (39)
Now we want to transform this into the cylinder coordinate $(t,r,\phi)$, so that we can match the coefficients in the expansion of BTZ and eventually solve the DT equations mode by mode. The transformation $(x(r,\phi),\sigma(r,\phi),t)$ is complex in the bulk, but simple near the boundary $r\to\infty$ and $x\to$-$\infty$
$e^{x}\sim\frac{1}{2\kappa\cos\phi}r^{-1},\quad\sigma=-\phi$ (40)
Note that this is only half of the entire conformal boundary. On the other half boundary $x\to\infty$, the boundary condition is the standard Dirichlet boundary condition for both Janus and BTZ with the discrete eigenvalues and eigenfunctions. In addition, since we only care about half of the range of $\phi$ at one side of the boundary, we use the orthogonal basis in $\phi\in[-\tfrac{\pi}{2},\tfrac{\pi}{2}]$. It is $B=\{\cos(2m\phi)\,\cup\,\sin((2m+1)\phi)\}_{m\in\mathbb{N}_{\geq 0}}$. (Although we demand that near the defect we recover the normalizable condition for the solutions, we need the complete orthogonal eigenfunction basis for both Neumann and Dirichlet boundary conditions due to the cos factor out front). We will expand both the BTZ and Janus solutions in this basis.
Let us write down the Janus wavefunction near the boundary (38) in terms of the basis $B$. It is convenient in that the even-parities $g_{\omega,k=even}$ are expanded in terms of $\cos(2m\phi)$, while the odd-parities $g_{\omega,k=odd}$ are expanded in terms of $\sin((2m+1)\phi)$. In addition, we have to add a rescaling factor $J_{m}\in\mathbb{R}$ for each basis
$\begin{split}\Psi_{\omega}^{Janus}\sim\sum_{m\geq 0}\Bigg[&\cos(2m\phi)\left(r^{-(2-\Delta_{J})}\frac{1}{(2\kappa)^{2-\Delta_{J}}}\sum_{k\geq 0}J_{2k}\,\alpha_{2k}^{J}\,f_{m,2k}^{\alpha}+r^{-\Delta_{J}}\frac{1}{(2\kappa)^{\Delta_{J}}}\sum_{k\geq 0}J_{2k}\,\beta_{2k}^{J}\,f_{m,2k}^{\beta}\right)\\ +\,&\sin((2m+1)\phi)\bigg(r^{-(2-\Delta_{J})}\frac{1}{(2\kappa)^{2-\Delta_{J}}}\sum_{k\geq 0}J_{2k+1}\,\alpha_{2k+1}^{J}\,f_{m,2k+1}^{\alpha}\\ &\quad+r^{-\Delta_{J}}\frac{1}{(2\kappa)^{\Delta_{J}}}\sum_{k\geq 0}J_{2k+1}\,\beta_{2k+1}^{J}\,f_{m,2k+1}^{\beta}\bigg)\Bigg]\end{split}$ (41)
Here
[…]
Notice that for $f_{0,2k}$ the coefficient in front of the integral will be $1/\pi$.
On the other hand, the expansion for quasinormal modes in BTZ black hole in (31) is simpler in this basis, with $\alpha_{m}^{BTZ}=\alpha_{-m}^{BTZ}$ and $\beta_{m}^{BTZ}=\beta_{-m}^{BTZ}$.
5.2 Coupling over half space
We first study the case where the entire half of the boundary CFT on one side of the interface is coupled to the bath via DT deformation. In the gravity, the coefficients of the sources are free, and we need to recover the freedom of rescaling the bath’s coefficients $\alpha_{m}^{BTZ}$ and $\beta_{m}^{BTZ}$ to a constant $b_{m}$, for each mode independently Karch and Youssef (2025). The reality condition together with the boundary condition that the solution vanishes at the boundaries $\phi=\pm\tfrac{\pi}{2}$ demand $b_{2m}=b_{-2m}:=B_{2m}/2\in\mathbb{R}$ and $b_{2m+1}=-b_{-2m-1}:=-i\,B_{2m+1}/2\in i\,\mathbb{R}$. This means that for scalar fluctuations in BTZ with dimension $\Delta_{BTZ}$, the expansion of the solution can be written as
$\begin{split}\Psi_{\omega}^{BTZ}&\sim\sum_{m\in\mathbb{Z}}b_{m}(\alpha_{m}^{BTZ}\,r^{-(2-\Delta_{BTZ})}+\beta_{m}^{BTZ}\,r^{-\Delta_{BTZ}})e^{im\phi}\\ &\sim\sum_{m\geq 0}\big[B_{2m}\cos(2m\phi)\left(\alpha_{2m}^{BTZ}\,r^{-(2-\Delta_{BTZ})}+\beta_{2m}^{BTZ}\,r^{-\Delta_{BTZ}}\right)\\ &\qquad\quad+B_{2m+1}\sin((2m+1)\phi)\left(\alpha_{2m+1}^{BTZ}\,r^{-(2-\Delta_{BTZ})}+\beta_{2m+1}^{BTZ}\,r^{-\Delta_{BTZ}}\right)\big]\end{split}$ (43)
In a marginal DT deformation, the dimension of the scalar in Janus and BTZ are complements $\Delta_{J}+\Delta_{BTZ}=2$. Below we use $\Delta:=\Delta_{BTZ}$ for simplicity in notations. We solve for $\{J_{k\geq 0},B_{m\geq 0},\omega\}$.
[…]
The cutoff for integer $k$ in $g_{\omega,k}$ and $J_{k}$ is $0\leq k\leq N_{\omega}:=\lfloor\mathrm{Re}(\omega)-\Delta_{Janus}\rfloor$ in order to keep the integral and functions convergent. We impose this same cutoff on the integer $m$ labeling the plane-wave basis in BTZ. This matches the number of DOF with the number of equations.
5.3 Source on a single point
If we consider coupling the BTZ and Janus on one single point ($\phi=\phi_{0}$) on one side instead of coupling the two on an entire half boundary, the setup is closer to Barad et al. (2025) where they studied relaxation across the interface for fermionic spin chains with dissipation introduced at a single site. In this case, we demand the source for each $e^{im\phi}$ mode be the same as it is a delta function source on the $\phi$ space. Using the relations in (42) and converting the basis to $\cos(2m\phi),\sin(2m+1)\phi$, we will solve for the equations below:
$\begin{split}B_{0}\alpha_{0}^{BTZ}(\omega,\Delta)&=\frac{B}{2},\\ B_{2m}\alpha_{2m}^{BTZ}(\omega,\Delta)&=B\cos(2m\phi_{0}),\\ B_{2m+1}\alpha_{2m+1}^{BTZ}(\omega,\Delta)&=B\sin((2m+1)\phi_{0}),\\ \frac{1}{(2\kappa)^{\Delta}}\sum_{k\geq 0}J_{2k}\,\alpha_{2k}^{J}(\omega,2-\Delta)\,f_{0,2k}^{\alpha}(\omega,2-\Delta)&=\frac{J}{2},\\ \frac{1}{(2\kappa)^{\Delta}}\sum_{k\geq 0}J_{2k}\,\alpha_{2k}^{J}(\omega,2-\Delta)\,f_{m,2k}^{\alpha}(\omega,2-\Delta)&=J\cos(2m\phi_{0}),\\ \frac{1}{(2\kappa)^{\Delta}}\sum_{k\geq 0}J_{2k+1}\,\alpha_{2k+1}^{J}(\omega,2-\Delta)\,f_{m,2k+1}^{\alpha}(\omega,2-\Delta)&=J\sin((2m+1)\phi_{0}),\\ \frac{J}{2}&=B_{0}h(2\Delta-2)\beta_{0}^{BTZ}(\omega,\Delta),\\ \frac{B}{2}=-h(2\Delta-2)\frac{1}{(2\kappa)^{2-\Delta}}\sum_{k\geq 0}J_{2k}&\,\beta_{2k}^{J}(\omega,2-\Delta)\,f_{0,2k}^{\beta}(\omega,2-\Delta)\end{split}$ (45)
6 Numerical results for quasinormal modes and $c_{\text{relax}}$
In this section we present the numerical analysis of the relaxation spectrum obtained by coupling the Janus geometry to a BTZ bath through the double-trace deformation described in Section 5. Throughout this section we fix the BTZ horizon radius to $r_{+}=\sqrt{0.8}$. We focus on the lowest-lying dissipative mode, which governs the late-time relaxation dynamics. In the numerical calculations carried out by Mathematica, we observed that the half-space coupling in section 5.2 is less stable than the setup in 5.3 where the coupling of the two systems is through a delta function source located on one side of the interface. Therefore, we focus on the delta-function coupling below and leave a detailed analysis of the half-space coupling to the future.
6.1 Liouvillian gap
For each value of the Janus parameter $\kappa$, we extract the quasinormal mode with the smallest absolute imaginary part,
$\Gamma_{\text{Janus}}(\kappa)\equiv-\text{min}\,[\mathrm{Im}\,\omega(\kappa)]\,$, (46)
which can be identified as the Liouvillian gap of the dual interface CFT, characterizing the relaxation time of the system Žnidarič (2015); Baumgartner and Narnhofer (2008).
To isolate the effect of the interface on the dissipation from one side to the other, we normalize this quantity by the corresponding decay rate in the planar AdS3 limit
$\Gamma_{\text{AdS}_{3}}\equiv-\text{min}\,[\mathrm{Im}\,\omega(\kappa=1)]\,.$ (47)
Note that for each fixed parameter of the system, there are an infinite tower of quasinormal modes. The branch with the smallest $\mathrm{Im}\,\omega$ is not necessarily the one with small $\mathrm{Re}\,w$. In fact, in our numerical results below, it is often the case that $\mathrm{Im}\,\omega$ decays to a constant when $\mathrm{Re}\,w$ approaches infinity in the tower. This constant is exactly the minimal imaginary part we desired. Below, the real part of the quasinormal modes $\mathrm{Re}\,w\approx 100$.
We then define the dimensionless relaxation ratio with and without the interface as a function Barad et al. (2025)
$c_{\text{relax}}\;\equiv\;\frac{\Gamma_{\text{Janus}}(\kappa)}{\Gamma_{\text{AdS}_{3}}(\kappa=1)}(\Delta,h,\phi_{0})\,.$ (48)
By construction, $c_{\text{relax}}=1$ in presence of a topological interface, and is 0 when the interface is totally reflective.
Physically, this ratio measures the suppression of the interface relaxation rate relative to the thermal dissipation without the interface, in analogy with equilibrium interface observables such as $c_{LR}$ or $c_{\text{eff}}$ quantifying energy or information transmission rate across the interface.
6.2 Universality of dissipation
As mentioned above, we consider the setup in which the Janus system is coupled to the BTZ bath through a localized source at angular position $\phi_{0}$ on the boundary, as described in Section 5.2. We first test its dependence on the angular position. As shown in Figure 1, up to our numerical precision, we found that the lowest decay rate is insensitive to the location of the coupling point:
$c_{\text{relax}}(\kappa,\Delta,h,\phi_{0})\equiv c_{\text{relax}}(\kappa,\Delta,h)\quad\text{for all }\phi_{0}\in\left(-\tfrac{\pi}{2},\tfrac{\pi}{2}\right)\,.$ (49)
Figure 1: The quasinormal modes of BPS-Janus for $\Delta=1.3$, $\kappa=1$ (blue) and $\kappa=1.5$ (red), at fixed coupling $g\equiv|(2\Delta-2)|h=0.1$. The green line is their ratio $c_{\text{relax}}$ shown on the right y-axis, signaling its independence of the position of insertion of dissipation.
The independence of the insertion point of the coupling supports our claim that the dissipation across the interface is a genuine interface observable that does not depend on the details of the excitation. Below we fix $\phi_{0}=0$.
Next we study the quasinormal modes at different double-trace coupling strengths $h$, at different conformal dimensions $\Delta$ of the scalar. For each of these parameters, we again solve for Eq. (45) for the lowest imaginary part of the quasinormal modes with and without the interface, as shown in Figure 2(a). In particular, at the conformally coupling points $\Delta=0.5,1.5$, there is a rigid transition at around $g\equiv|(2\Delta-2)|h=1$ where the modes switch between an AdS3 branch and a BTZ branch, signaling the strong/weak transition raised in Karch and Youssef (2025).
The numerical results also suggest that at fixed Janus deformation $\kappa$, the ratio $c_{\text{relax}}$ is independent of both $h$ and $\Delta$, as shown in Figure 2.
$c_{\text{relax}}(\kappa,\Delta,h)\equiv c_{\text{relax}}(\kappa)\quad\text{for all }\,h\,,\,\tfrac{1}{2}<\Delta<\tfrac{3}{2}\,$, (50)
(a) Quasinormal modes and their ratio $c_{\text{relax}}$ (green, right y-axis) at $\kappa=1.2,\,\Delta=1.4$ versus double-trace coupling $g$
(b) Quasinormal modes and $c_{\text{relax}}$ at $\kappa=1.2,\,g=0.5$ versus the conformal dimension $\Delta$ within the unitarity bound.
Figure 2: The dependence of $c_{\text{relax}}$ on $h$ and $\Delta$ is within our numerical precision ($10^{-4}$), supporting its universality.
6.3 $c_{\text{relax}}$
Figure 3: $c_{\text{relax}}$ (red dots) versus the BPS-Janus deformation parameter $\kappa$. The green line is the energy transmission coefficient $c_{LR}=1/\kappa^{2}$, and the blue line is the entanglement entropy across the interface $c_{\text{eff}}=1/\kappa$, which almost exactly fits $c_{\text{relax}}$ up to our numerical precision.
Finally, we examine the universal function $c_{\text{relax}}(\kappa)$ on the BPS-Janus interface parameter and compare it with other interface observables like the energy transmission coefficient $c_{LR}$ Baig et al. (2024) and the entanglement entropy of half space $c_{\text{eff}}$ Gutperle and Miller (2016).
$c_{LR}=\frac{1}{\kappa^{2}}<c_{\text{relax}}(\kappa)\;\approx\;\frac{1}{\kappa}=c_{\text{eff}},$ (51)
as illustrated in Figure 3. This is different from the example of the spin chain studied in Barad et al. (2025), where $c_{\text{relax}}$ was the lowest of the three interfaces central charges.
Our results provide strong evidence that $c_{\text{relax}}$ is a universal quantity characterizing dissipative dynamics across conformal interfaces, insensitive to details of the coupling to the bath and governed primarily by the interface geometry.
7 Conclusion and future directions
Our main result was to demonstrate that within one particular holographic example, the supersymmetric Janus geometry coupled to a finite temperature bath, the relaxation of dissipation across an interface is completely universal in the same way it was for the previously considered case of spin chains: the ratio of the dissipation rate between the case with and without interface is independent of the source of dissipation, most notably its location, its strength and its microscopic realization. While the rates themselves depend on these details, the ratio does not and defines a new central-charge like quantity that is universal to the interface. While the previously identified $c_{\text{eff}}$ and $c_{LR}$ were interpreted as transmission coefficients for information and energy respectively, $c_{\text{relax}}$ similarly can be interpreted as a transmission coefficient for dissipation.
Given that this universality now has been seen in two very different examples one may strongly suspect that this is in fact true more generally, but it would be good to work out a few more cases to check, such as the half-space coupling in Section 5.2. Beyond that, it would be important to understand how the values of the 3 quantities related to each other. It was previously argued that $c_{LR}$ is always less than $c_{\text{eff}}$ Karch et al. (2024a). However, when it comes to $c_{\text{relax}}$ the situation is quite different. While it was found in the spin chain that $c_{\text{relax}}$ was always below the other two, in our case we find that $c_{\text{relax}}$ is larger than $c_{LR}$ and, within the limits of our precision, identical to $c_{\text{eff}}$. This latter identity may be an artifact of our supersymmetric setup.
Acknowledgements
We’d like to thank Xueda Wen and Merna Youssef for useful discussions. This work was supported in part by DOE grant DE-SC0022021 and by a grant from the Simons Foundation (Grant 651678, AK).